"What are the missing numbers? This first of two numbers is 9 more than the second. Three times the first number is equal to 7 more than 5 times the second."
Answer by stanbon(75887)

"What are the missing numbers?
The first of two numbers is 9 more than the second.
Three times the first number is equal to 7 more than 5 times the second."
------------
1st: x
2nd: y
---
Equations:
x = y+9
3x= 5y+7
---------------
Substitute for "x" and solve for "y":
3(y+9) = 5y+7
3y+27 = 5y+7
2y = 20
y = 10
------
Since x = y+9, x = 19
============================
Cheers,
Stan H.